OpenStudy (alyssacheeseburger):

Decide whether the equation is a trigonometric identitiy. Explain your reasoning sin^2theta csc^2theta =sin^2theta +cos^2theta

OpenStudy (johnweldon1993):

I would, first, write everything in terms of \(\large sin\) and \(\large cos\)

OpenStudy (alyssacheeseburger):

and how would i do that haha I'm completely lost sorry

OpenStudy (johnweldon1993):

Don't worry about it lol Well we have \[\large sin^2\theta \times csc^2\theta = sin^2\theta + cos^2\theta\] We want to have EVERYTHING be in terms of sin and cos...so if I tell you that \[\large csc^2\theta = \frac{1}{sin^2\theta}\] We would have \[\large sin^2\theta \times \frac{1}{sin^2\theta} = sin^2\theta + cos^2\theta\] And what would be next?

OpenStudy (alyssacheeseburger):

You can cancel Sin^2theta and 1/sin^2theta ?

OpenStudy (johnweldon1993):

Exactly \[\large \frac{sin^2\theta}{sin^2\theta} = sin^2\theta + cos^2\theta\] Meaning \[\large 1 = sin^2\theta + cos^2\theta\] And on the right side...what do we know about \(\large sin^2\theta + cos^2\theta\) ?

OpenStudy (alyssacheeseburger):

do they equal one?

OpenStudy (johnweldon1993):

Lol yes they do, that is one of the FIRST identities to know is \(\large sin^2x + cos^2x = 1\) So we have \(\large 1 = 1\) and therefor we have proven it is an identity

OpenStudy (alyssacheeseburger):

thank you so much haha as you can tell I'm falling behind in class

OpenStudy (johnweldon1993):

Ehh it happens, just don't try and rush into whats going on NOW...make sure you understand these basics first :)